OpenStudy (anonymous):

How would u graph these system of inequalities? 20x + 15y ≤ 950

OpenStudy (anonymous):

this is only one inequalities, if it's a syestem then there should at leat another inequalities

OpenStudy (anonymous):

x >_ 20 y >_ 15

OpenStudy (anonymous):

20x + 15y ≤ 950 x >_ 20 y >_ 15

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

let first 20x+15y=950 take x=1--->y=930/15 =62 then take x=4---->y=870/15 =58 plot on the x-y plane these two points to form a straight line

OpenStudy (anonymous):

draw the vertical line x=-20

OpenStudy (anonymous):

draw the horizontal line y=-15

OpenStudy (anonymous):

the solution would be the area bounded by the three lines

OpenStudy (anonymous):

so wait what are the 3 points? 1, 62 4, 58 -20, -15?

OpenStudy (anonymous):

two points for the first line (1,62),(4,48)

OpenStudy (anonymous):

at (-20,0) draw a vertical line that intersct the x axis only

OpenStudy (anonymous):

at (0,-15) draw a horz. line that interstects with y axis only

OpenStudy (anonymous):

awesome thanks

OpenStudy (anonymous):

yw